# Network-Realized Model Predictive Control Part II: Distributed Constraint Management

### v1

@@ 2. $b2)$
$\mathbf{K}_{\mathbf{D}i}(z)=\left[\scriptsize\begin{array}[]{c|c}A_{wi}-zI_{n_{wi}}&B_{wi}\\ \hline\cr C_{wi}&D_{wi}\end{array}\right]:=\left[\tiny\begin{array}[]{ccc|c}A_{r(\alpha_{ui}+1)}-zI_{n_{r(\alpha_{ui}+1)}}&&&B_{r(\alpha_{ui}+1)}\\ &\ddots&&\vdots\\ &&A_{r(\alpha_{ui}+n_{ui})}-zI_{n_{r(\alpha_{ui}+n_{ui})}}&B_{r(\alpha_{ui}+n_{ui})}\\ \hline\cr C_{r(\alpha_{ui}+1)}&&&O\\ &\ddots&&\vdots\\ &&C_{r(\alpha_{ui}+n_{ui})}&O\\ \end{array}\right].\vspace{1mm}$ (8)
3.2 Interfacing with the First Layer
Our control architecture’s first layer is implemented as depicted in Figure 2, in which the distributed control laws are represented by the TFM
$\mathbf{K}_{\mathbf{D}}(z):=\begin{bmatrix}\mathbf{K}_{\mathbf{D}1}^{\top}(z)&\dots&\mathbf{K}_{\mathbf{D}N}^{\top}(z)\end{bmatrix}^{\top},\vspace{-3mm}$ (8)
where each $\mathbf{K}_{\mathbf{D}i}(z)=S_{ui}^{\top}\mathbf{K}_{\mathbf{D}}(z)$ is the TFM of the $i^{\text{th}}$ area’s local subcontroller. Crucially, these subcontrollers are implemented via the states-space realizations given in (8), located at the top of the next page, whose component matrices (see Section 4.1 in [10] for more details) are given, for all $\ell\in\mathcal{A}_{ui}$, by
$\left\{\begin{array}[]{ll}\widetilde{A}_{r\ell}:=\tiny\begin{bmatrix}-a_{1\ell}&\dots&-a_{(n_{r\ell}-1)\ell}\end{bmatrix}^{\top},&A_{r\ell}:=\ \tiny\begin{bmatrix}\widetilde{A}_{r\ell}&I_{n_{r\ell}-1}\\ -a_{n_{r\ell}\ell}&O\end{bmatrix},\\ B_{r\ell}:=\tiny\begin{bmatrix}K_{1i}^{\top}&\dots&K_{n_{r\ell}i}^{\top}\end{bmatrix}^{\top},&C_{r\ell}:=\tiny\ \ \begin{bmatrix}1&0&\dots&0\end{bmatrix}.\end{array}\right.$ (10)
Henceforth, we define the orders $n_{wi}:=\textstyle\sum_{\ell=1}^{n_{ui}}n_{r\ell}$ along with $n_{w}:=\textstyle\sum_{i=1}^{N}n_{wi}$. We also denote the state vectors of the implementations from (8) as $w_{i}[k]$ and their initial conditions as $w_{ci}\in\mathbb{R}^{n_{wi}}$. Moreover, we concatenate the latter into $w_{c}:=\tiny\begin{bmatrix}w_{c1}^{\top}&\dots&w_{cN}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{n_{w}}$, which are the initial conditions of the first layer’s global implementation.
@@ 3.2 Interfacing with the First Layer
$u_{fi}[k]:=S_{{ui}}^{\top}u_{f}[k]=\mathbf{K}_{\mathbf{D}i}(z)\star\tiny\begin{bmatrix}u_{f}[k]+\beta_{f}[k]\\ x[k]+\zeta[k]+u_{s1}[k]+\beta_{s1}[k]\end{bmatrix},\normalsize$ (11)
for all $i\in\{1:N\}$, while employing only information from each area’s neighbourhood. Algebraically, this reduces to the fact that, for all $j\in\{1:N\}\backslash\mathcal{N}_{i}$, the implementations from (8)-(10) satisfy
$\mathrm{col}_{q}(B_{r\ell})=O,\,\forall\,\ell\in\mathcal{A}_{ui},\,q\in\mathcal{A}_{xj}\cup(\{n_{x}\}\oplus\mathcal{A}_{uj}).\vspace{-3mm}$ (12)
@@
Although we consider the first layer’s control laws to be obtained through an NRF-based framework, as in our companion paper [10], note that any set of distributed subcontrollers implemented as in (8)-(12) may be employed for the procedures described in the sequel.
The other main feature of the NRF-based layer is revealed by first introducing the exogenous signal vector
$d_{s}[k]:=\begin{bmatrix}(\zeta[k]+\beta_{s1}[k])^{\top}&\beta_{s2}^{\top}[k]&\beta_{f}^{\top}[k]&d^{\top}[k]\end{bmatrix}^{\top},\vspace{-3mm}$ (13)
in which $\zeta[k]$ represents state measurement noise, $\beta_{f}[k]$ stands for the first layer’s communication disturbance, $\beta_{s1}[k]$ and $\beta_{s2}[k]$ denote encoding errors and communication disturbance associated with the second layer’s command signals and $d[k]$ is the disturbance associated directly with the networks’ dynamics in (4a).
@@
We assume that the communication infrastructure of our control architecture is endowed error-detecting/correcting transmission mechanisms. Therefore, all communication errors affecting transmitted information are modelled as disturbance which originates at the source of transmission, not on the receiving ends.
By using the signal vector introduced in (13), while also merging the MPC subcontroller’s command signals from Figure 1 into the vectors $u_{s1}[k]:=\tiny\begin{bmatrix}u_{s11}^{\top}[k]&\dots&u_{s1N}^{\top}[k]\end{bmatrix}^{\top}$ and $u_{s2}[k]:=\tiny\begin{bmatrix}u_{s21}^{\top}[k]&\dots&u_{s2N}^{\top}[k]\end{bmatrix}^{\top}$, it is possible to express the first layer’s closed-loop dynamics as follows
$\scriptsize\begin{bmatrix}x[k]\\ u_{f}[k]\end{bmatrix}=\mathbf{F}_{\mathbf{Q}}(z)\scriptsize\begin{bmatrix}I_{(n_{x}+n_{u})}\\ O\end{bmatrix}\star\scriptsize\begin{bmatrix}u_{s1}[k]\\ u_{s2}[k]\end{bmatrix}+\\ +\mathbf{F}_{\mathbf{Q}}(z)\star d_{s}[k]+\mathcal{I}_{\mathbf{Q}}[k]{\scriptsize\begin{bmatrix}x_{c}\\ w_{c}\end{bmatrix}}\,,$ (14)
where $\mathbf{F}_{\mathbf{Q}}(z)$ is a strictly proper ($\lim_{z\rightarrow\infty}\mathbf{F}_{\mathbf{Q}}(z)=O$) TFM which models the closed-loop’s response to exogenous inputs, and $\mathcal{I}_{\mathbf{Q}}[k]$ is a time-varying matrix which governs the closed-loop system’s evolution with respect to initial conditions. The main result of our companion paper [10] shows that $\mathbf{F}_{\mathbf{Q}}(z)$ and $\mathcal{I}_{\mathbf{Q}}[k]$ are easily tunable via a freely chosen $\mathbf{Q}$-parameter (hence, the employed notation), thus enabling the complete design of the first layer’s closed-loop response. It is with respect to the dynamics given in (14) that we formulate both the objectives and the design procedure of our second layer.
3.3 Objectives of the MPC Layer
Given the high-level functioning principles described in Section 3.1 and the closed-loop dynamics summarized in Section 3.2, our MPC-based second layer must simulta-
neously achieve the following objectives:
  1. $\mathrm{I})$
    it employs tractable and area-based prediction models, which completely capture the local behaviour of the dynamics given in (14) (Section 4);
  2. $\mathrm{II})$
    it formulates and manages all of its constraints in a distributed manner, by adapting set-theoretical techniques [3] in accordance with the area partitioning from (5)-(6) (Section 5);
  3. $\mathrm{III})$
    it admits a formulation based upon receding horizon optimization (see, for example, Chapter 1 in [7]) which regulates only local variables $\{x_{i},u_{fi}\}$, but ensures global constraint satisfaction and recursive feasibility (Section 6);
The remainder of this paper is dedicated to presenting the solutions to these design challenges.
4 Obtaining an MPC-Oriented Formulation
@@ 4.1 Nominal Area-based Prediction Models
In order to formulate tractable prediction models for our architecture’s second layer, we need to account for the state variables of the implementations from (8). Thus, aside from the $x$ and $u$ subscripts appearing in (6)-(7), we introduce the additional subscript $w$, in order to refer to the state vector of each area’s NRF subcontroller, and we extend the index collection defined originally in (5) to $\mathcal{A}_{i}:=(\mathcal{A}_{xi},\mathcal{A}_{ui},\mathcal{A}_{wi}),\,\forall\,i\in\{1:N\}$.
We begin our reformulation of the first-layer’s closed-loop dynamics by pointing out that, due to the fact that the TFM $\mathbf{F}_{\mathbf{Q}}(z)$ from (14) is strictly proper, it is always possible to obtain state-space realizations of type
@@
for all $i\in\{1:N\}$. By defining now $Z_{i}:=\mathrm{diag}(S_{xi},S_{ui})$ along with ${Z}_{ci}:=\mathrm{diag}(S_{xi},S_{wi})$, we employ the realizations from (15) to describe the MPC-oriented area model of the $i^{\text{th}}$ area, which can be expressed as
@@
    $\xi_{i}[k]$ is the state vector of the realization obtained in (15) and which always has zero initial conditions at initial time $k_{0}$ (recall the identity from (3));
@@
$\left\{\begin{split}\psi_{i}[k]:=&\ Z_{i}^{\top}\mathbf{F}_{\mathbf{Q}}(z)\star d_{s}[k],\\ \psi_{xi}[k]:=&\ \tiny\begin{bmatrix}I_{n_{xi}}&O\end{bmatrix}\psi_{i}[k],\ \psi_{ui}[k]:=\ \tiny\begin{bmatrix}O&I_{n_{ui}}\end{bmatrix}\psi_{i}[k],\end{split}\right.\vspace{-3mm}$ (17)
@@
$\left\{\begin{split}&\theta_{i}[k]:=\textstyle\sum_{j\in\mathcal{N}_{i}}Z_{i}^{\top}\mathcal{I}_{\mathbf{Q}}[k]Z_{cj}\tiny\begin{bmatrix}x_{cj}\\ w_{cj}\end{bmatrix},\normalsize\\ &\theta_{xi}[k]:=\tiny\begin{bmatrix}I_{n_{xi}}&O\end{bmatrix}\theta_{i}[k],\ \theta_{ui}[k]:=\tiny\begin{bmatrix}O&I_{n_{ui}}\end{bmatrix}\theta_{i}[k],\end{split}\right.\vspace{-3mm}$ (18)
@@
$\left\{\begin{split}&\delta_{i}[k]:=\sum_{j\in\{1:N\}\backslash\{i\}}Z_{i}^{\top}\mathbf{F}_{\mathbf{Q}}(z)\tiny\begin{bmatrix}Z_{j}\\ O\end{bmatrix}\star\begin{bmatrix}u_{s1j}[k]\\ u_{s2j}[k]\end{bmatrix}+\\ &\qquad\qquad\quad\,+\sum_{j\in\{1:N\}\backslash\mathcal{N}_{i}}Z_{i}^{\top}\mathcal{I}_{\mathbf{Q}}[k]Z_{cj}\tiny\begin{bmatrix}x_{cj}\\ w_{cj}\end{bmatrix},\\ &\delta_{xi}[k]:=\tiny\begin{bmatrix}I_{n_{xi}}&O\end{bmatrix}\delta_{i}[k],\ \delta_{ui}[k]:=\tiny\begin{bmatrix}O&I_{n_{ui}}\end{bmatrix}\delta_{i}[k],\end{split}\right.\vspace{-3mm}\normalsize$ (19)
@@ 1. $i)$
    available to a second layer subcontroller, represented by $\theta_{xi}[k]$ and $\theta_{ui}[k]$;
@@ 2. $ii)$
    unavailable to a second layer subcontroller, represented by $(\psi_{xi}[k]+\delta_{xi}[k])$ and $(\psi_{ui}[k]+\delta_{ui}[k])$.
@@
The quantities described in (19) are called residual due to the fact that it is the primary goal of the first layer to attenuate the effect that these signals have on the area dynamics from (16a)-(16d). Indeed, since these terms represent the contribution of all the variables which are not available (in terms of measurement or prediction) to the $i^{\text{th}}$ area, the only sensible recourse is to minimize their influence (see Section 5 in [10]).
@@ 4.2 Accounting for Inaccurate Initial Conditions
$\tiny\begin{bmatrix}\widetilde{x}_{c}\\ \widetilde{w}_{c}\end{bmatrix}:=\tiny\begin{bmatrix}{x}_{c}+\nu_{x}\\ {w}_{c}+\nu_{w}\end{bmatrix},\quad\tiny\begin{bmatrix}\widetilde{x}_{c{i}}\\ \widetilde{w}_{ci}\end{bmatrix}:=\tiny\begin{bmatrix}{x}_{ci}+S^{\top}_{{xi}}\nu_{x}\\ {w}_{ci}+S^{\top}_{{wi}}\nu_{w}\end{bmatrix},\normalsize\vspace{-3mm}$ (20)
for any index $i\in\{1:N\}$ along with any $\tiny\begin{bmatrix}\nu_{x}^{\top}\ \,\nu_{w}^{\top}\end{bmatrix}^{\top}\in\mathcal{V}$.
@@
$\left\{\begin{split}\widetilde{\theta}_{i}[k]:=&\ \theta_{i}[k]+\sum_{j\in\mathcal{N}_{i}}Z_{i}^{\top}\mathcal{I}_{\mathbf{Q}}[k]Z_{cj}\tiny\begin{bmatrix}S^{\top}_{{xj}}\nu_{x}\\ S^{\top}_{{wj}}\nu_{w}\end{bmatrix},\\ \widetilde{\theta}_{xi}[k]:=&\ \tiny\begin{bmatrix}I_{n_{xi}}&O\end{bmatrix}\widetilde{\theta}_{i}[k],\ \widetilde{\theta}_{ui}[k]:=\ \tiny\begin{bmatrix}O&I_{n_{ui}}\end{bmatrix}\widetilde{\theta}_{i}[k],\end{split}\right.\vspace{-3mm}$ (22)
@@
$\left\{\begin{split}\widetilde{\delta}_{i}[k]:=&\ \delta_{i}[k]+\sum_{j\in\{1:N\}\backslash\mathcal{N}_{i}}Z_{i}^{\top}\mathcal{I}_{\mathbf{Q}}[k]Z_{cj}\tiny\begin{bmatrix}S^{\top}_{{xj}}\nu_{x}\\ S^{\top}_{{wj}}\nu_{w}\end{bmatrix},\\ \widetilde{\delta}_{xi}[k]:=&\ \tiny\begin{bmatrix}I_{n_{xi}}&O\end{bmatrix}\widetilde{\delta}_{i}[k],\ \widetilde{\delta}_{ui}[k]:=\ \tiny\begin{bmatrix}O&I_{n_{ui}}\end{bmatrix}\widetilde{\delta}_{i}[k],\end{split}\right.\vspace{-3mm}\normalsize$ (23)
@@ 4.3 Connecting the Augmented Prediction Models to the Nominal Ones
$x_{i}[k]=\widetilde{x}_{i}[k]-\eta_{xi}[k],\quad u_{fi}[k]=\widetilde{u}_{fi}[k]-\eta_{ui}[k],\vspace{-3mm}$ (24)
@@
$\left\{\begin{split}\eta_{i}[k]:=&\ \textstyle\sum_{j=1}^{N}Z_{i}^{\top}\mathcal{I}_{\mathbf{Q}}[k]Z_{cj}\tiny\begin{bmatrix}S^{\top}_{{xj}}\nu_{x}\\ S^{\top}_{{wj}}\nu_{w}\end{bmatrix},\\ \eta_{xi}[k]:=&\ \tiny\begin{bmatrix}I_{n_{xi}}&O\end{bmatrix}\eta_{i}[k],\ \eta_{ui}[k]:=\ \tiny\begin{bmatrix}O&I_{n_{ui}}\end{bmatrix}\eta_{i}[k].\end{split}\right.\normalsize\vspace{-3mm}$ (25)
Having obtained the area-based prediction models in (16a)-(16d) and having augmented them as in (21a)-(21b), it follows that objective $I)$ from Section 3.3 has been fully addressed. We now proceed to discuss constraint formulation in the distributed setting, and the means by which the identities from (24) can be employed to ensure constraint satisfaction (in the presence of unreliable initial conditions) for the variables in (16b)-(16c).
@@ 5 Constraints in the Distributed Setting
The first step in breaking down the second layer’s distributed synthesis problem into locally solvable subproblems is to assign independent control objectives to each of the network’s $N$ areas. Therefore, if the goal is to ensure that the signal vectors $x$ and $u$ are constrained (for all instances of time, relative to an initial moment $k_{0}$) to a pair of sets $\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$ and $\mathcal{U}\subseteq\mathbb{R}^{n_{u}}$, respectively, then our procedure must begin by defining the pairs of sets $(\mathcal{X}_{i},\mathcal{U}_{i})$, with $i\in\{1:N\}$, such that
$\left\{\begin{split}x[k]\in\mathcal{X}&\Longleftrightarrow x_{i}[k]\in\mathcal{X}_{i},\,\forall\,i\in\{1:N\},\\ u[k]\in\mathcal{U}&\Longleftrightarrow u_{i}[k]\in\mathcal{U}_{i},\,\forall\,i\in\{1:N\}.\end{split}\right.\vspace{-3mm}$ (26)
@@
The assumption that the pair $(\mathcal{X},\mathcal{U})$ can be decomposed as in (26) is not only required, when attempting to break down the global problem into any configuration of local subproblems, but it is also a mild one (see its equivalent counterpart in [1]-[2]). Indeed, the conditions located on the left-hand side of (26) are often represented (in practice) by box constraints, for which the aforementioned decomposition may always be performed.
The aim of the subcontrollers which form the second layer (recall Figure 1) is to ensure the global, left-hand side inclusions from (26) by independently satisfying the local ones, on the right-hand side of (26). With respect to the components of the closed-loop dynamics from (14), this problem can be broken down into five subproblems:
@@ 2. $ii)$
    translating the command-related objectives from (26) into first layer command constraints;
@@ 5. $v)$
We now proceed to tackle each of these subproblems, in the order in which they are indicated.

### v4

@@ 4.1 Interfacing with the First Layer
[…]
for all $i\in\{1:N\}$, while employing only information from each area’s neighbourhood. Algebraically, this reduces to the fact that, for all $j\in\{1:N\}\setminus\,\mathcal{N}_{i}$, the implementations from (10)-(11) satisfy
$\mathrm{col}_{q}(B_{r\ell})=O,\,\forall\,\ell\in\mathcal{A}_{ui},\,q\in\mathcal{A}_{xj}\cup(\{n_{x}\}\oplus\mathcal{A}_{uj}).\vskip-8.53581pt$ (13)
@@
Although we consider the first layer’s control laws to be obtained through an NRF-based framework, as in our companion paper [part1], note that any set of distributed subcontrollers implemented akin to (9)-(13) may be employed for the procedures described in the sequel.
By using the signal vector introduced in (7), it is possible to express the first layer’s closed-loop dynamics as
$\scriptsize\begin{bmatrix}x[k]\\ u_{f}[k]\end{bmatrix}=\left(\mathbf{F}_{\mathbf{Q}}(z)\scriptsize\begin{bmatrix}I_{(n_{x}+n_{u})}\\ O\end{bmatrix}\right)\star\scriptsize\begin{bmatrix}u_{s1}[k]\\ u_{s2}[k]\end{bmatrix}+\\ +\mathbf{F}_{\mathbf{Q}}(z)\star d_{s}[k]+\mathcal{I}_{\mathbf{Q}}[k]{\scriptsize\begin{bmatrix}x_{c}\\ w_{c}\end{bmatrix}}\,$, (14)
where $\mathbf{F}_{\mathbf{Q}}(z)$ is a strictly proper TFM, which is to say that $\lim_{z\rightarrow\infty}\mathbf{F}_{\mathbf{Q}}(z)=O$. Notably, $\mathbf{F}_{\mathbf{Q}}(z)$ models the closed-loop’s response to exogenous inputs, while $\mathcal{I}_{\mathbf{Q}}[k]$ is a time-varying matrix which governs the closed-loop system’s evolution with respect to its initial conditions.
@@
While the signal vectors $d[k]$ and $\zeta[k]$ in (7) and (implicitly) in (14) possess classical interpretations in system-theoretical literature, the three vectors denoted generically as $\beta_{\bullet}[k]$ in (7) are caused primarily by encoding errors in the employed communication infrastructure. Due to finite-precision representation, the absolute values of the signals in $\beta_{\bullet}[k]$ may be considered to be upper bounded and, unless stated otherwise, the samples making up $\beta_{\bullet}[k]$ are implicitly assumed to be uniformly distributed within their respective bounds. In addition to this, we point out that boundedness (not necessarily in terms of amplitude) also extends to both $d[k]$ and $\zeta[k]$, in most practical applications. Indeed, due to the stabilising properties of NRF-based control, the effect of all these signals upon the closed-loop system can be made bounded in some sense, by an appropriate choice of systemic norm optimisation (see Chapter 4 of [zhou] and Section 5.2 of [part1]). Moreover, the results from [PUB] can be neatly incorporated into our considered approach, if one wishes to delve further into the stochastic setting.
The main result of our companion paper [part1] shows that $\mathbf{F}_{\mathbf{Q}}(z)$ and $\mathcal{I}_{\mathbf{Q}}[k]$ are easily tunable via a freely chosen $\mathbf{Q}$-parameter (hence, the employed notation), thus enabling the complete design of the first layer’s closed-loop response. The full details pertaining to the synthesis and tuning of these closed-loop maps are given in our companion paper [part1], yet, for the benefit of the reader, we have included a summary of the key properties ensured by NRF-based control in Appendix A.
@@ 4.2 Nominal Area-based Prediction Models
In order to formulate tractable prediction models for our architecture’s second layer, we need to account for the state variables of the implementations from (10). Thus, aside from the $x$ and $u$ subscripts appearing in (5)-(6), we introduce the additional subscript $w$ to refer to the state vector of each area’s NRF subcontroller, and we extend the index collection defined originally in (4) to $\mathcal{A}_{i}:=(\mathcal{A}_{xi},\mathcal{A}_{ui},\mathcal{A}_{wi}),\,\forall\,i\in\{1:N\}$.
We begin our reformulation of the first layer’s closed-loop dynamics by pointing out the fact that, since the TFM $\mathbf{F}_{\mathbf{Q}}(z)$ from (14) is strictly proper, it is always possible to obtain state-space realisations of type
@@
for all $i\in\{1:N\}$. By defining now $Z_{i}:=\mathrm{diag}(S_{xi},S_{ui})$ along with ${Z}_{ci}:=\mathrm{diag}(S_{xi},S_{wi})$, we employ the realisations from (15) to describe the MPC-oriented area model of the $i^{\text{th}}$ area, which can be expressed as
@@
    $\xi_{i}[k]$ is the state vector of the realisation obtained in (15) and which always has zero initial conditions at initial time $k_{0}$ (recall the identity from (3));
@@
[…]
@@
[…]
@@
[…]
@@
    available to a second-layer subcontroller, represented by $\theta_{xi}[k]$ and $\theta_{ui}[k]$;
@@
    unavailable to a second-layer subcontroller, represented by $(\psi_{xi}[k]+\delta_{xi}[k])$ and $(\psi_{ui}[k]+\delta_{ui}[k])$.
@@
The quantities described in (19) are called residual because it is the primary goal of the first layer to attenuate the effect that these signals have on the area dynamics from (16a)-(16d). Since these terms, along with those in (17), represent the contribution of variables which are not available (in terms of measurement or prediction) to the $i^{\text{th}}$ area, the only sensible recourse is to minimise their influence (see Section 5 in [part1]).
@@ 4.3 Accounting for Inaccurate Initial Conditions
[…]
for any index $i\in\{1:N\}$, along with any $\tiny\begin{bmatrix}\nu_{x}^{\top}\ \,\nu_{w}^{\top}\end{bmatrix}^{\top}\in\mathcal{V}$.
@@
[…]
@@
[…]
@@ 4.4 Connecting the Augmented Prediction Models to the Nominal Ones
$x_{i}[k]=\widetilde{x}_{i}[k]-\eta_{xi}[k],\quad u_{fi}[k]=\widetilde{u}_{fi}[k]-\eta_{ui}[k],\vskip-8.53581pt$ (24)
@@
[…]
Having obtained the area-based prediction models in (16a)-(16d) and having augmented them as in (21a)-(21b), we now proceed to discuss constraint formulation for the optimisation problems described in (8).
@@ 5 Constraints in the Distributed Setting
The most important step in breaking down the second layer’s distributed synthesis problem into locally solvable subproblems is to assign independent control objectives to each of the network’s $N$ areas. Recalling the sets $\mathcal{X}$ and $\mathcal{U}$, introduced in Section 3.2, we begin by defining the pairs of sets $(\mathcal{X}_{i},\mathcal{U}_{i})$, with $i\in\{1:N\}$, which satisfy
[…]
@@
The assumption that the pair $(\mathcal{X},\mathcal{U})$ can be decomposed as in (26) is not only required, when attempting to break down the global problem into any configuration of local subproblems, but it is also a mild one (see its equivalent counterpart in [DMPC1, DMPC2]). Indeed, the conditions located on the left-hand side of (26) are often represented (in practice) by box constraints, for which said decomposition may always be performed.
The chief aim of the subcontrollers which form the second layer (recall (8) and Figure 1) is to ensure the global, left-hand side inclusions from (26) by independently satisfying the local ones, on the right-hand side of (26). With respect to the closed-loop dynamics from (14), this problem can be broken down into five components:
@@
    translating the command-related objectives from (26) into first-layer command constraints;
@@
We now proceed to tackle each of these subproblems, in the order in which they are indicated above.